OpenStudy (spring98):

If (7^2)^p = 7^6, what is the value of p?

OpenStudy (spring98):

does p equal 3?

OpenStudy (spring98):

@zzr0ck3r

OpenStudy (spring98):

can anyone plz help me?

OpenStudy (spring98):

@Loser66 @pooja195 @mathlover2014 @NightCrawlers_

OpenStudy (freckles):

yes 2 time 3=6

OpenStudy (freckles):

\[(7^2)^p=7^6 \\ \text{ which is equivalent to } 7^{2p}=7^6 \text{ which means that } 2p=6 \] and yes 2*3=6 so p=3

OpenStudy (spring98):

thanks!!!

OpenStudy (freckles):

np